Nico is saving money for his college education. He invests some money at 8​%, and ​$1900 less than that amount at 3 %. The investments produced a total of ​$240 interest in 1 yr. How much did he invest at each​ rate?
Found 2 solutions by addingup, fractalier:
Answer by addingup(3677)

.08x+.03(x-1900) = 240
.08x+.03x-57 = 240
0.11x = 297
x = 2700 is the amount invested at 8% and
2700-1900 = 800 is invested at 3%
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2700*.08 = 216
800* .03 = 24
__________--------
Total . . . 240 we have the correct answer

Answer by fractalier(6550)

Call how much Nico invests at 8%, x.
Then the amount invested at 3% must be x - 1900.
The setup looks like this
.08x + .03(x - 1900) = 240
We then distribute and solve for x...
.08x + .03x - 57 = 240
.11x - 57 = 240
.11x = 297
Now divide by .11 and get
x = 297 / .11 = $2700 at 8%
and thus $1900 less means
$800 at 3%...